# A General-Purpose and Unified Framework for

### v2

@@ Abstract
Note: Trinity-RFT is currently under active development. This technical report (arXiv v2 version) corresponds to commit id 63d4920 (July 14, 2025) of the GitHub repository, and will be continuously updated as the codebase evolves. Comments, suggestions and contributions are welcome!
@@ 2.1.1 Unified Support for Diverse RFT Modes
We conclude this subsection with two remarks. (1) Given the unified implementation of various RFT modes, it is easy to design and implement a hybrid mode with Trinity-RFT that combines multiple modes into a single learning process. One example is learning with both online rollout data and offline-collected expert data, via jointly optimizing two loss terms corresponding to these two data sources. Section 3.2 illustrates how to implement this conveniently in Trinity-RFT. (2) We take a system-algorithm co-design perspective in the development of Trinity-RFT, aiming to unify and generalize diverse RFT methodologies in this framework. RFT-core provides the necessary infrastructure for achieving this goal. While the main text of this technical report focuses on the system perspective, interested readers may refer to Appendix A for further discussions on the algorithm perspective of off-policy/asynchronous RL.
@@ 3.3.2 Experiments: Real Learning with Vanilla GRPO
We aim to compare the real learning processes by different RL modes. For simplicity and controlled variability, we use the vanilla GRPO [28] algorithm for all RL modes, without specific algorithm design for asynchronous or off-policy cases. GRPO mainly relies on the mechanism of clipping probability ratio (defaults to the range of ) to handle the off-policyness of experiences. For future works, we will investigate more advanced off-policy or asynchronous RL algorithms (e.g., the OPMD algorithm proposed by Kimi k1.5 [34] and two OPMD variants developed in Appendix A), and develop new ones to accommodate diverse RL modes. To encourage the exploration of the rollout model, we disable the Kullback-Leibler (KL) penalty or loss in our experiments.
@@ 4 Conclusion and Next Steps
Further development of Trinity-RFT includes enhancing support for generic multi-step agentic workflows, incorporating more advanced RL algorithms (especially off-policy or asynchronous ones), making the choices of hyperparameters more adaptive and less reliant on manual tuning, augmenting data pipelines with smarter sampling strategies and data processing operations, and conducting more thorough experiments and evaluations with Trinity-RFT.
@@ Acknowledgements
Appendix A Off-Policy Reinforcement Learning Algorithms
For off-policy learning, one might simply adopt offline methods like supervised fine-tuning and direct preference optimization (DPO) [24], or their iterative variants where (some of) the training samples are generated online by the policy model itself and filtered by their reward values, e.g., rejection-sampling fine-tuning and iterative DPO. There also exist more advanced RL algorithms with dedicated design for off-policy or asynchronous scenarios [20, 25, 34, 6, 22, 41, 44, 45].
In the rest of this section, we present some of our findings about off-policy RL (with a focus on the bandit setting) through the lens of online policy mirror descent (OPMD). At the end of this section, we arrive at the surprising conclusion that the standard policy gradient, weighted by a coefficient and using the group mean reward as the baseline, can be a feasible direction for updating the policy even in off-policy settings (while the standard theory of policy gradient only holds for on-policy settings). This has been validated empirically in our exploratory experiments during the development of Trinity-RFT 141414https://github.com/modelscope/Trinity-RFT/blob/main/docs/sphinx_doc/source/tutorial/example_reasoning_advanced.md.
A.1 OPMD: Kimi’s Version
Below is a recap of the OPMD variant proposed in the technical report of Kimi k1.5 [34].
Analysis.
For a specific task/query and a reference policy , consider the following objective for training policy at a particular iteration of the RL process:
Note that can be changing during the RL process. In [34], is set to at the -th iteration, i.e., when updating the policy from to .
The optimal policy for this objective satisfies the following: for any response ,
Taking logarithm of both sides of Eq. (1), we see that the optimal policy must satisfy the following consistency condition:
Algorithm.
Based on the above analysis, [34] proposes the following OPMD variant. For a query , first sample rollouts from the reference policy, then define a surrogate loss as follows:
Although this is an off-policy method (since the rollout policy is different from the policy being updated), it is still limited because the rollouts have to be sampled from the particular policy for the -th iteration of the RL process, as mentioned earlier. The reason for this limitation is the need of estimating using samples from .
A.2 Pairwise OPMD
Analysis.
To eliminate the term, we note that Eq. (1) is equivalent to the following:
Taking logarithm of both sides, we have
or equivalently,
Note that this holds true for a pair of arbitrary responses and .
Algorithm.
For a query and arbitrary responses , we define the following surrogate loss:
Here can be any reference policy for KL regularization, regardless of how were sampled. While this is a fully off-policy RL method, it has its own limitation: to run this algorithm, we should make sure that multiple (at least 2) rollouts for the same task are included within one micro-batch (whose size is typically much smaller that of a batch or mini-batch), which adds to infrastructure complexity.
Remark 1.
In the special case of , the above method, termed as “pairwise OPMD”, turns out to be the same as “contrastive policy gradient” proposed in [6], albeit with a simpler and more intuitive derivation.
A.3 OPMD: an Embarrassingly Simple Variant
Analysis.
Consider the -th iteration of the RL process, i.e., updating from to , and use as the reference policy. For a specific task/query , recall from Section A.1 the original objective:
We leverage the analysis in Section A.2, and take a closer look at the following pairwise loss for and , normalized by to make the loss scale invariant to the hyperparameter :
The trick here is that, if we only intend to take one gradient step of this loss at , then the value of is simply zero. As a result,
and thus
where in the last line.
Algorithm.
To this end, we update from to by taking one gradient step of the following surrogate loss, where we simplify the constant factor from to and also drop the factor:
This is exactly the standard policy gradient using the group mean reward as the baseline, but derived differently with an off-policy interpretation. The hyperparameter controls the size of each policy update. A larger is effectively the same as a smaller learning rate.
Remark 2.
As a heuristic, when additional regularization with respect to a fixed policy (e.g., a SFT model ) is desired, we simply add a regularization term (denoted by ) to the above objective:

### v3

@@ Abstract
Note: Trinity-RFT is currently under active development. This technical report corresponds to commit id 63d4920 (July 14, 2025) of the GitHub repository, and will be continuously updated as the codebase evolves. Comments, suggestions and contributions are welcome!
@@ 2.1.1 Unified Support for Diverse RFT Modes
We conclude this subsection with two remarks. (1) Given the unified implementation of various RFT modes, it is easy to design and implement a hybrid mode with Trinity-RFT that combines multiple modes into a single learning process. One example is learning with both online rollout data and offline-collected expert data, via jointly optimizing two loss terms corresponding to these two data sources. Section 3.2 illustrates how to implement this conveniently in Trinity-RFT. (2) We take a system-algorithm co-design perspective in the development of Trinity-RFT, aiming to unify and generalize diverse RFT methodologies in this framework. RFT-core provides the necessary infrastructure for achieving this goal. This technical report focuses on the system perspective, and we refer interested readers to the literature for recent algorithmic developments in off-policy / asynchronous RL for LLMs [21, 6, 26, 35, 7, 23, 42, 45, 46, 47].
@@ 3.3.2 Experiments: Real Learning with Vanilla GRPO
We aim to compare the real learning processes by different RL modes. For simplicity and controlled variability, we use the vanilla GRPO [29] algorithm for all RL modes, without specific algorithm design for asynchronous or off-policy cases. GRPO mainly relies on the mechanism of clipping probability ratio (defaults to the range of ) to handle the off-policyness of experiences. For future works, we will investigate more advanced off-policy or asynchronous RL algorithms, and develop new ones to accommodate diverse RL modes. To encourage the exploration of the rollout model, we disable the Kullback-Leibler (KL) penalty or loss in our experiments.
@@ 4 Conclusion and Next Steps
Further development of Trinity-RFT includes incorporating more advanced RL algorithms (especially off-policy or asynchronous ones), making the choices of hyperparameters more adaptive and less reliant on manual tuning, augmenting data pipelines with smarter sampling strategies and data processing operations, and conducting more thorough experiments and evaluations with Trinity-RFT.